# An Objective Improvement Approach

### v3

@@ 2. Preliminaries
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution [^3] If $\sigma$ is not co-optimal, then there is a maximiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}(v)<w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ or there is a minimiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}(v)>w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$. In both cases, the inequation for this particular edge $e$ is violated..
@@ 2.1. Simplex method
A common way to compute $\mathsf{val}_{\mathcal{G}}$ is to create a linear programming instance from $\mathcal{G}$. In this paragraph, we provide the background for solving a linear programming instance and outline the Simplex method that is the standard algorithm used to find a solution.
@@
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that maximises (or minimises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Given a game $\mathcal{G}$, we can define an objective function represented by the vector c, the input matrix $A$, and the vector $b$, such that the solution $x$ to the linear programming instance corresponds to $\mathsf{val}_{\mathcal{G}}$, as follows. In this translation from $\mathcal{G}$ to the linear programming instance, the matrix $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a nonzero value (with the exception of self-loops that have only one nonzero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}\cdot x_{j}+a_{ik}\cdot x_{k}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$ and $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), with $a_{ij}=1$, $a_{ik}=\lambda_{e}$, and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative sign for the minimiser, and the objective function maximises the sum of the values of $x$. Every linear programming instance can be expressed in a simpler standard form by applying transformations to the objective function and to the constraints. A standard form requires to find $x$ that minimises $c\cdot x$ such that $B\cdot x=b$ and $x\geq 0$, with $b$ positive and $B$ a sub-matrix of $n$ linear independent rows vectors of $A$. Such a form can always be obtained by introducing new variables, called slack, that transform the inequations into equations. Note that none of the algorithms for solving DPGs requires standard form instances (as described in the next section, the strategy improvement approach solves single-player DPGs with a fixed objective function, while the approach we present in this work solves two-players DPGs but on a different objective function at each iteration), however, to describe the Simplex method, in this section we assume the linear programming instance is translated in standard form.
From a graphical point of view, the set of costraints, i.e., the rows of $B$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope is a solution. The objective function, however, provides an additional vector that shows the direction of better solutions. As a result, if there is an optimal solution (and there always is an optimal solution in the linear programming problems we consider in this paper), then there is an optimal solution that lies on a corner of the polytope. This observation allows to restrict the search space for optimal solutions to the corners of the polytope. In fact, it suffices to move along the edges of the polytope to find an optimal solution.
From an algebraic point of view, each non-zero variable in a solution $x$ is active. The subset of $n$ constraints of $A$ selected by $B$ is called basis, since every constraint of $A$ can be written as a unique combination of the $n$ constraints in $B$. If two bases $B$ or solutions $x$ differ for one active variable, then they are adjacent.
The Simplex method solves instances in standard form by finding a solution and then by moving to an adjacent solution that improves the value of the objective function. In the graphical representation, two adjacent solutions lie on adjacent corners. The algorithm stops when the adjacent solutions do not improve the value of the objective function of the one currently selected.
input : The elements of a linear programming problem $A,b,c$
output : The optimal solution vector $x$
1 $B\leftarrow{}$ random initial basis matrix from $A$
2 while true do
    foreach $j$ non-basis index do
       3 $c^{\prime}_{j}\leftarrow{}$ reduced cost over objective function for index j
    end foreach
   4if $c^{\prime}\geq 0$ then
       return $x$
    end if
   5Select $j$ such that $c^{\prime}_{j}<0$
   6 Select $k$ such that $k$ is the min ratio test index
   7 $B\leftarrow{}\mathsf{Pivot}(A,b,c,j,k)$
   8 $x\leftarrow{}\mathsf{Update}(B,c)$
end while
Algorithm 1 Simplex method
The procedure is reported in Algorithm 1. At Line 1, an initialisation step selects a first basis (a subset of $n$ rows from $A$ which are linearly independent, thus $B$ is a square $n$-matrix). The “for each” loop at Line 3 computes, for every index $j$ of a variable that does not belong (i.e., not active) to the current basis $B$, the difference between the current objective function value and the one obtained by selecting a different variable $j$. Such a difference is called reduced cost. In case all these differences are non-negative, that is, when the vector $c^{\prime}$ composed of the $c^{\prime}_{j}$ costs is non-negative, then the current solution cannot be improved and it is returned at Line 4. The index $j$ corresponds to a column of $B$ and identifies the candidate variable to enter the basis. Otherwise, at Line 5, one variable from the current basis is selected to be replaced by $j$. The leaving variable $k$ is found by computing for each constraint (row of $B$) the ratio between the constant value from $b$ corresponding to the current row and the coefficient of the candidate variable to join the basis. If the $k$-th row has such a minimum value, then one of the improving adjacent corners is selected as solution by replacing $k$ with $j$. This check is called minimum ratio test, while the move from a basis to an adjacent one is called pivoting operation (Line 7) and consists in column operations over $B$ that maintain the linear independence of the matrix. Once the basis matrix has been changed, the solution vertex can be updated (Line 8).
@@ 3. Outline and Motivation Example
If the maximiser decides to stay forever in his vertex (using the self-loop), his overall reward in the play that starts at (and, due to his choice, always stays in) his vertex, is 00. If he decides to move on to the minimiser vertex the $n^{th}$-time, then the reward is $\frac{\lambda^{n}}{1-\lambda}$.
The optimal decision of the maximiser is, therefore, to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}(v)=w_{e}+\lambda_{e}\mathsf{val}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
@@
The unique valuation that satisfies these inequations and produces a sharp inequation (i.e., satisfied as equation) for some outgoing edge of each vertex assigns $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$. This valuation also defines the optimal strategies of the players (to stay for the minimiser, and to move on for the maximiser).
Solving a discounted payoff game means finding this valuation and/or these strategies.
@@
This term is always non-negative, since it corresponds to the sum of the edges’ contributions that are all non-negative. Moreover, when only optimal strategies are selected to form this objective function, the value 00 can be taken, and where it is taken, it defines the correct valuation.
As the maximiser’s choice to take the self-loop is not optimal, the resulting objective function the strategies define, that is $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$, cannot reach 00. But let us take a look at what an optimal solution w.r.t. this objective function looks like.
@@
Thus, in this toy example, while selecting the wrong edge cannot result in the objective function taking the value 00, we still found the optimal solution. In general, we might need to update the objective function. To update the objective function, we change the outgoing edges of some (or all) vertices so that the overall value of the objective function goes down. Note that this can be done not only when the linear program returns an optimal solution but also during its computation. For example, when using a simplex method, updating the objective function can be used as an alternative pivoting rule at any point during the traversal of the simplex.
In general, the valuation returned as solution is computed by using objective functions based on strategies that are not necessarily optimal. We now adjust the example so that choosing both self-loops is not only non-optimal, but also does not define the correct valuation. To this end, we use different discount factors [^4] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1: we choose $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three resulting inequations are (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$; (2) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(b)$; and (3) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(a)$. Making the adjusted inequations (2) and (3) sharp still results in the values $\mathsf{val}(a)=\mathsf{val}(b)=0$, and the objective function still takes the value of $1$. While making inequations (1) and (3) sharp provides the values $\mathsf{val}(a)=3$ and $\mathsf{val}(b)=2$; the objective function takes the value $\frac{4}{3}$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp still conflicts with inequation (3).
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f)$
These steps can be identified in Algorithm 2. The objective function is fixed at the beginning (Line 1) and corresponds to the sum of the values of all the vertices. The algorithm simply maximises it [^5] The optimal minimiser strategy chooses edge that minimises the value for every minimiser vertex. In the linear program, this is revlected by an inequeation for every outgoing edge, like the inequation $\mathsf{val}(a)\leq 1+\mathsf{val}(a)$ from the running example, and then maximises the value, which forces the selected value to be the minimal value such that one such inequation is sharp. by searching for a solution that maximises the value of all vertices. (1) The strategy of one of the two players (for convention the maximiser – if we want to instead fix the strategy of the minimiser, we would have to minimise the sum of the values for all vertices instead) is initialised at Line 2. (2) Before the computation of the valuation at Line 5, the set of constraints is updated accordingly to the selected strategy of the maximiser (Line 4). At this point, if the strategy is optimal w.r.t. the current valuation, then the algorithm ends. Otherwise, (3) there exists an improvement, which is applied at Line 7.
@@
Thus, in our novel symmetric approach, the constraints remain while the objective is updated; in strategy improvement, instead, the objective remains while the constraints are updated. Moreover, the players and their strategies are treated quite differently in strategy improvement algorithms: while the candidate strategy of the maximiser results in making the inequations of the selected edges sharp (and dropping all other inequations of maximiser edges), the optimal counter strategy is found by minimising the objective. This is again in contrast to our novel symmetric approach, which treats both players equally.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
This objective function $f_{\sigma}$ is given to a linear programming algorithm, alongside with the inequations set $H$. We underline that, due to the inequation $I_{(v,v^{\prime})}$, the value of $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))$ is non-negative for all $(v,v^{\prime})\in E$ in any valuation $\mathsf{val}$ (optimal or not) that satisfies the system of inequations $H$.
@@
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, according to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We refer to this unique set of values as the evaluation of $\mathbf{b}$, denote by $\mathsf{val}_{\mathbf{b}}$.
The call $\mathsf{LinearProgramming}(H,f_{\sigma})$ to some linear programming algorithm returns a valuation $\mathsf{val}$ of the vertices that minimises $f_{\sigma}$ while satisfying $H$ (Line 5); for convenience, we require this valuation to also be the evaluation $\mathsf{val}_{\mathbf{b}}$ for some basis $\mathbf{b}$ of $H$. (Note that the simplex algorithm, for example, only uses valuations of this form in every step.) We call this valuation a valuation associated to $\sigma$.
We say that a valuation $\mathsf{val}$ defines strategies of both players if, for every vertex $v\in V$, the inequation of (at least) one of the outgoing edges of $v$ is sharp. These are the strategies defined by using, for every vertex $v\in V$, an outgoing edge for which the inequation is sharp. Note that there can be more than one of these inequations for some of the vertices.
@@
While we discuss how to implement this key function in the next section, we observe here that the algorithm terminates with a correct result with any implementation that chooses a better objective function in each round: correctness is due to it only terminating when $\mathsf{val}$ defines strategies for both players, which implies (cf. Theorem 1) that $\mathsf{val}$ is the valuation of $\mathcal{G}$ ($\mathsf{val}=\mathsf{val}_{\mathcal{G}}$) and all strategies defined by $\mathsf{val}$ are co-optimal. Termination is obtained by a finite number of positional strategies: by Observation 4, the value of the objective function of all of them is non-negative, while the objective function of an optimal solution to co-optimal strategies is 00 (cf. Theorem 1), which meets the termination condition of Line 6 (cf. Observation 4).

### v4

@@ 2. Preliminaries
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution [^3] If $\sigma$ is not co-optimal, then there is a maximiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}_{\sigma}(v)<w_{e}+\lambda_{e}\mathsf{val}_{\sigma}(v^{\prime})$ or there is a minimiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}_{\sigma}(v)>w_{e}+\lambda_{e}\mathsf{val}_{\sigma}(v^{\prime})$. In both cases, the inequation for this particular edge $e$ is violated..
@@ 2.1. Simplex method
A common way to compute $\mathsf{val}_{\mathcal{G}}$ is to create a sequence of linear programming instances from $\mathcal{G}$. In this paragraph, we provide the background for solving a linear programming instance and outline the Simplex method that is the standard algorithm used to find a solution.
@@
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that minimises (or maximises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Given a game $\mathcal{G}$, we can define an objective function represented by the vector c, the input matrix $A$, and the vector $b$, such that the solution $x$ to the linear programming instance corresponds to $\mathsf{val}_{\mathcal{G}}$, as follows. In this translation from $\mathcal{G}$ to the linear programming instance, the matrix $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a nonzero value (with the exception of self-loops that have only one nonzero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}\cdot x_{j}+a_{ik}\cdot x_{k}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$ and $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), with $a_{ij}=1$, $a_{ik}=-\lambda_{e}$, and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative sign for the minimiser, and the objective function maximises the sum of the values of $x$.
As we use no details of any implementation of a simplex algorithm, we only rehash the principle way it operates here.
From a graphical point of view, the set of constraints, i.e., the rows of $A$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope (including its surface) is a solution.
Broadly speaking, the simplex algorithm traverses the corner of the polytope so that the value the linear objective function takes for the corner is improving – in our case decreasing as we minimise. More precisely, it picks subsets of $n$ inequations that are made sharp: when red as equations, they define a valuation in $\mathbb{R}^{n}$ that satisfies all $m$ inequations; they are bases and define the corners of the simplex.
For this, the simplex algorithm operates in two phases. In a first phase it finds any basis that defines such a corner. We will not discuss this phase as it is not relevant for this article [^4] It is also the case that, in all but the first call to a linear program, we already have a basis to start from..
In the second phase, the simplex algorithm essentially updates the basis to a neighbouring one, meaning that only one inequation is removed from and one inequation is entered into the set of sharp inequations. When making this pick, the simplex algorithm only allows for such updates that are (1) still bases (i.e. their rows in $A$ must be linearly independent to that they define a point in $\mathbb{R}^{n}$); (2) the point in $\mathbb{R}^{n}$ they define satisfies all $m$ inequations, and (3) the value of the objective function improves.
For non-optimal solutions (w.r.t. the objective function), this is always possible, though in some cases the improvement in (3) is not strict. Even in those cases, there is still a series of changes that all satisfy (1-3) to an optimal solution. Those cases are called degenerate, and while we do not discuss how they are treated in simplex algorithms, we point out that the problems and their solutions are similar to what we describe in Sections 5.2 and 5.3, respectively.
@@ 3. Outline and Motivation Example
If the maximiser decides to stay forever in his vertex (using the self-loop), his overall reward in the play that starts at (and, due to his choice, always stays in) his vertex, is 0. If he decides to move on to the minimiser vertex the $n^{th}$-time, then the reward is $\frac{\lambda^{n}}{1-\lambda}$.
The optimal decision of the maximiser is, therefore, to move on the first time, which yields the maximal reward of $\frac{\lambda}{1-\lambda}$. Every vertex $v$ has some outgoing edge(s) $e=(v,v^{\prime})$ where $\mathsf{val}_{\mathcal{G}}(v)=w_{e}+\lambda_{e}\mathsf{val}_{\mathcal{G}}(v^{\prime})$ holds [ZP96]; these edges correspond to the optimal decisions for the respective player.
@@
The unique solution that satisfies these inequations and produces a sharp inequation (i.e. satisfied as equation) for some outgoing edge of each vertex assigns $\mathsf{val}(a)=\frac{1}{1-\lambda}$ and $\mathsf{val}(b)=\frac{\lambda}{1-\lambda}$. This solution also defines the optimal strategies of the players (to stay for the minimiser, and to move on for the maximiser).
Solving a discounted payoff game means finding this solution and/or these strategies.
@@
This term is always non-negative, since it corresponds to the sum of the edges’ contributions that are all non-negative. Moreover, when only optimal strategies are selected to form this objective function, the value 0 can be taken, and where it is taken, it defines the correct valuation of the game.
As the maximiser’s choice to take the self-loop is not optimal, the resulting objective function the strategies define, that is $1-(1-\lambda)\mathsf{val}(a)+(1-\lambda)\mathsf{val}(b)$, cannot reach 0. But let us take a look at what an optimal solution w.r.t. this objective function looks like.
@@
Thus, in this toy example, while selecting the wrong edge cannot result in the objective function taking the value 0, we still found the optimal solution. In general, we might need to update the objective function. To update the objective function, we change the outgoing edges of some (or all) vertices so that the overall value of the objective function goes down. Note that this can be done not only when the linear program returns an optimal solution but also during its computation. For example, when using a simplex method, updating the objective function can be used as an alternative pivoting rule at any point during the traversal of the simplex.
In general, the valuation returned as solution is computed by using objective functions based on strategies that are not necessarily optimal. We now adjust the example so that choosing both self-loops is not only non-optimal, but also does not define the correct valuation of the DPG. To this end, we use different discount factors [^5] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1: we choose $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three resulting inequations are (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$; (2) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(b)$; and (3) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(a)$. Choosing both self loops results in the objective function to minimise $1-(\frac{1}{3})\mathsf{val}(a)+\frac{2}{3}\mathsf{val}(b)$.
Making the adjusted inequations (2) and (3) sharp still results in the values $\mathsf{val}(a)=\mathsf{val}(b)=0$, and the objective function still takes the value of $1$. While making inequations (1) and (3) sharp provides the values $\mathsf{val}(a)=3$ and $\mathsf{val}(b)=2$; the objective function takes the value $\frac{4}{3}$ at this point. Finally, if we consider the last combination, making (1) and (2) sharp still conflicts with inequation (3).
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f)$
These steps can be identified in Algorithm 1. The objective function is fixed at the beginning (Line 1) and corresponds to the sum of the values of all the vertices. The algorithm simply maximises it [^6] The optimal minimiser strategy chooses edge that minimises the value for every minimiser vertex. In the linear program, this is revlected by an inequeation for every outgoing edge, like the inequation $\mathsf{val}(a)\leq 1+\lambda\mathsf{val}(a)$ from the running example, and then maximises the value, which forces the selected value to be the minimal value such that one such inequation is sharp. by searching for a solution that maximises the value of all vertices. (1) The strategy of one of the two players (for convention the maximiser – if we want to instead fix the strategy of the minimiser, we would have to minimise the sum of the values for all vertices instead) is initialised at Line 2. (2) Before the computation of the valuation at Line 5, the set of constraints is updated accordingly to the selected strategy of the maximiser (Line 4). At this point, if the strategy is optimal w.r.t. the current valuation, then the algorithm ends. Otherwise, (3) there exists an improvement, which is applied at Line 7.
@@
Thus, in our novel symmetric approach, the constraints remain while the objective is updated; in strategy improvement, instead, the objective remains while the constraints are updated. Moreover, the players and their strategies are treated quite differently in strategy improvement algorithms: while the candidate strategy of the maximiser results in making the inequations of the selected edges sharp (and dropping all other inequations of maximiser edges), the optimal counter strategy is found by maximising the objective. This is again in contrast to our novel symmetric approach, which treats both players equally.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
This objective function $f_{\sigma}$ is given to a linear programming algorithm, alongside with the inequations set $H$. We underline that, due to the inequation $I_{(v,v^{\prime})}$, the value of $\mathsf{offset}(\mathsf{val},(v,v^{\prime}))$ is non-negative for all $(v,v^{\prime})\in E$ in any solution $\mathsf{val}$ (optimal or not) that satisfies the system of inequations $H$.
@@
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, according to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We refer to this unique set of values as the evaluation of $\mathbf{b}$, denote by $\mathsf{val}_{\mathbf{b}}$. Note that $\mathsf{val}_{\mathbf{b}}$ is in particular a solution of $H$, so that all all $|E|$ inequations are satisfied.
The call $\mathsf{LinearProgramming}(H,f_{\sigma})$ to some linear programming algorithm returns a solution $\mathsf{val}$ of the vertices that minimises $f_{\sigma}$ while satisfying $H$ (Line 5); for convenience, we require this solution to also be the evaluation $\mathsf{val}_{\mathbf{b}}$ for some basis $\mathbf{b}$ of $H$. (Note that the simplex algorithm, for example, only uses solutions of this form in every step.) We call this solution a solution associated to $\sigma$.
We say that a solution $\mathsf{val}$ defines strategies of both players if, for every vertex $v\in V$, the inequation of (at least) one of the outgoing edges of $v$ is sharp. These are the strategies defined by using, for every vertex $v\in V$, an outgoing edge for which the inequation is sharp. Note that there can be more than one of these inequations for some of the vertices.
@@
While we discuss how to implement this key function in the next section, we observe here that the algorithm terminates with a correct result with any implementation that chooses a better objective function in each round: correctness is due to it only terminating when $\mathsf{val}$ defines strategies for both players, which implies (cf. Theorem 1) that $\mathsf{val}$ is the valuation of $\mathcal{G}$ ($\mathsf{val}=\mathsf{val}_{\mathcal{G}}$) and all strategies defined by $\mathsf{val}$ are co-optimal. Termination is obtained by a finite number of positional strategies: by Observation 4, the value of the objective function of all of them is non-negative, while the objective function of an optimal solution to co-optimal strategies is 0 (cf. Theorem 1), which meets the termination condition of Line 6 (cf. Observation 4).